OpenStudy (anonymous):

if the complex number Z lies on the circle |Z-2i| = 2sqrt(2) then find arg( (Z-2)/Z+2) ) ?

OpenStudy (anonymous):

Omg.. Hate complex no. !

mathslover (mathslover):

Unfortunately, I love complex no. :( :)

OpenStudy (anonymous):

Carry on bro ! :D D:

mathslover (mathslover):

Did I say that I know how to solve this question? :P , please wait, I am trying

mathslover (mathslover):

@Yahoo! try to put : z = x + iy

mathslover (mathslover):

then continue.

OpenStudy (anonymous):

|Z-2i| = 2sqrt(2) is a circle with centre (0,2) and radius 2sqrt(2)///

OpenStudy (anonymous):

and i think we have to do something with that..:)

mathslover (mathslover):

I can say that : \(x^2 - y^2 + 4y = 12\)

mathslover (mathslover):

\(x+iy = Z\)

OpenStudy (anonymous):



OpenStudy (anonymous):

hey guys ho is these character

OpenStudy (anonymous):

LoL It's you @jamespatrick1071

OpenStudy (anonymous):

U r Ryt @phi but i dont get u

OpenStudy (anonymous):

\[|Z-2i| = 2\sqrt{2}\]  \[arg( \frac{Z-2}{Z+2} )=arg(z-2)-arg(z+2)\]